# MoLA: Motion Generation and Editing with Latent Diffusion Enhanced by Adversarial Training

### v3

@@ Abstract.
In text-to-motion generation, controllability as well as generation quality and speed has become increasingly critical. The controllability challenges include generating a motion of a length that matches the given textual description and editing the generated motions according to control signals, such as the start-end positions and the pelvis trajectory. In this paper, we propose MoLA, which provides fast, high-quality, variable-length motion generation and can also deal with multiple editing tasks in a single framework. Our approach revisits the motion representation used as inputs and outputs in the model, incorporating an activation variable to enable variable-length motion generation. Additionally, we integrate a variational autoencoder and a latent diffusion model, further enhanced through adversarial training, to achieve high-quality and fast generation. Moreover, we apply a training-free guided generation framework to achieve various editing tasks with motion control inputs. We quantitatively show the effectiveness of adversarial learning in text-to-motion generation, and demonstrate the applicability of our editing framework to multiple editing tasks in the motion domain. Our code is available at https://github.com/sony/MoLA
Motion generation, Latent diffusion model, Adversarial training, Guided generation
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: XXXjournalvolume: XXXjournalnumber: XXXarticle: XXXccs: Computing methodologies Computer visionccs: Computing methodologies Motion processing
@@
Figure 1. MoLA achieves fast and high-quality human motion generation given textual descriptions while enabling motion editing applications. With MoLA, we can deal with various types of motion editing tasks in a single framework.
1. Introduction
@@
Human motion synthesis from text is an emerging task with highly relevant applications in fields such as multimedia production and computer animation. For example, an animator might wish to create or edit a motion prototype to verify their artistic intent before time-consuming animation or motion capture commences, generate smooth in-between motion between two motion capture clips, or generate specific motion that follows a predefined trajectory. In Figure 1, we demonstrate results of our approach on such motion generation and editing tasks. To be useful in those real-world applications, a method has to excel in three domains: (1) motion quality, which encompasses both the general motion quality and the adherence to the textual description; (2) fast inference time; (3) easy motion editing.
@@
Several recent works have attempted to address these desired properties: Methods based on vector quantization (VQ), such as T2M-GPT (Zhang et al., 2023b), MoMask (Guo et al., 2024), MMM (Pinyoanuntapong et al., 2024b), ParCo (Zou et al., 2024), and BAMM (Pinyoanuntapong et al., 2024a), achieve impressive generation quality by compressing human motion into discrete tokens and then sampling those tokens to synthesize motion. Diffusion-based methods in data space, such as MDM (Tevet et al., 2023), provide impressive flexibility with regard to quality and motion editability. To further improve motion quality and inference time, latent-space based methods such as MLD (Chen et al., 2023) or MotionMamba (Zhang et al., 2025) operate in a learned continuous latent space.
@@
However, no state-of-the-art method excels in all three domains necessary for real-world applications. For example, while latent-space based approaches such as VQ-based methods or MLD achieve impressive motion quality and fast inference time, they cannot edit a given motion sequence in a training-free manner. In contrast, data-space based methods such as MDM are able to edit motion in a training-free manner, which, however, comes at the cost of slow inference time and lower generation quality. Moreover, while these models excel at motion generation, they require manual control of the generated motion length, which limits their flexibility. For instance, MoMask (Guo et al., 2024) uses a length estimator during inference to predict motion length based on text input. However, inaccurate predictions may result in significant motion drift (Wu et al., 2024).
@@
To close this gap, we propose MoLA, Motion Generation and Editing with Latent Diffusion Enhanced by Adversarial Training. We revisit the representation of motion features used as inputs and outputs of the model and introduce an activation variable that characterizes the length of the motion. We utilize a variational auto-encoder (VAE) (Kingma and Welling, 2013) for its continuous latent space, which allows training-free editing. This is in contrast to discrete latent spaces that do not allow for training-free motion editing. To bridge the motion quality gap, we enhance the VAE training with adversarial learning, which has been shown to work well in the image (Esser et al., 2021; Iashin and Rahtu, 2021; Rombach et al., 2022) and audio (Liu et al., 2023) domains. We empirically show that our model achieves variable-length motion generation aligned with textual descriptions and high generation performance on a commonly used dataset. We also demonstrate that training-free guided diffusion can enable multiple motion editing tasks such as path-following, in-betweening, and upper body editing. These experiments also highlight the significant improvement in speed and performance of our model compared to the performance of existing training-free motion editing models (see Figure 2).
@@
The main contributions of this paper are threefold. First, we introduce an activation variable into the motion representation and show that our method can perform variable-length motion generation conditioned on text. Second, we propose a new continuous latent-based motion generation model that introduces adversarial training into the motion VAE and quantitatively show that it significantly pushes the limits of existing continuous-based methods. Finally, we demonstrate that our model not only shows high generation performance but also can deal with various types of motion editing tasks (path following, motion in-betweening, and upper body editing) in a training-free manner.
@@
Figure 2. Comparison of inference cost, generation performance, and editability for text-to-motion methods. $\bullet$ means a method that can edit motion in a training-free manner, and $\times$ means a method that cannot edit motion in a training-free manner. We calculate the average inference time on the test set of HumanML3D without model or data loading parts. All tests are performed on the same NVIDIA A100 GPU. The pink arrow in the figure indicates that our method significantly extends the performance boundaries (in terms of generation quality and speed) of methods categorized as enabling training-free editing.
2. Related Work
2.1. Motion Generation
@@
Text-to-motion generation technology has made rapid progress with the diffusion models and VQ-based models. MDM (Tevet et al., 2023) and MotionDiffuse (Zhang et al., 2024) adopt diffusion models for motion generation, which leads to better performance in terms of generation quality. However, these methods directly apply diffusion processes to raw motion sequences, thus resulting in slow generation. MLD (Chen et al., 2023) mitigates this issue by adopting a diffusion model in a low-dimensional latent space provided by a VAE trained on motion data, inspired by latent diffusion models (Rombach et al., 2022). VQ-based models have been studied as well, inspired by the success of VQ-based models in image generation (Gu et al., 2022; Yu et al., 2022; Chang et al., 2023). In this approach, a VQ-VAE model is first trained on motion data to acquire discrete motion representations, and deep generative models are then applied to generate sequences of discrete representations (also called tokens). T2M-GPT (Zhang et al., 2023b), AttT2M (Zhong et al., 2023), MotionGPT (Jiang et al., 2023) and ParCo (Zou et al., 2024) utilize autoregressive (AR) models to generate motion tokens. However, AR models are slow in inference because motion tokens are generated sequentially. To address this issue, M2DM (Kong et al., 2023) and DiverseMotion (Lou et al., 2023) apply discrete diffusion models to motion tokens in a latent space, whereas MMM (Pinyoanuntapong et al., 2024b), MoMask (Guo et al., 2024) and BAMM (Pinyoanuntapong et al., 2024a) adopt a mask prediction model.
@@
Figure 3. The overall framework of MoLA. Stage 1: A motion VAE enhanced by adversarial training learns a low-dimensional latent representation of diverse motion sequences. Stage 2: A text-conditioned latent diffusion model leverages this representation for fast and high-quality text-to-motion generation. Guided generation: During inference, a gradient-based method minimizes a loss function $\mathfrak{L}_{\text{Motion}}$ for each desired editing task, enabling multiple motion editing tasks within a unified framework.
2.2. Motion Editing
@@
Motion editing has attracted much research interest as well. MDM (Tevet et al., 2023) demonstrated upper body editing and motion in-betweening by applying diffusion inpainting to motion data in both the spatial and temporal domains. LGD (Song et al., 2023) demonstrated path-following motion generation with a guided diffusion that utilizes multiple samples from a suitable distribution to reduce bias. GMD (Karunratanakul et al., 2023) guides the position of the root joint to control motion trajectories. OmniControl (Xie et al., 2024) controls any joints at any time by guiding a pre-trained motion diffusion model with an analytic function. DNO (Karunratanakul et al., 2024) optimizes the diffusion latent noise of a pre-trained text-to-motion model with user-provided criteria in the motion space and achieves multiple editing tasks. However, these methods employ data-space diffusion models similar to MDM and have not been demonstrated with a latent diffusion model. Recently, MMM (Pinyoanuntapong et al., 2024b) demonstrated motion editing by placing masked tokens in the place that needs editing and applying the mask prediction framework, and MotionLCM (Dai et al., 2024) proposed a fast controllable motion generation framework by introducing latent consistency distillation and the motion ControlNet (Zhang et al., 2023a) manipulation in the latent space. However, their methods cannot control an arbitrary body part in a training-free manner.
3. Method
@@
The goal of this study is to develop a framework for fast and high-quality text-guided motion generation and to deal with multiple control tasks in a training-free way. To achieve this, we propose the following training and inference tricks (I)-(IV): (I) We review the representation of motion features to achieve improved accuracy and variable length motion generation. We select the necessary features to reduce the burden on our model’s VAE encoder and add an activation variable to the representation to determine the motion length (Section 3.1). (II) We train a motion VAE enhanced by adversarial training to achieve high generation performance (Section 3.2). (III) To reduce computational complexity while simultaneously enabling high-quality text-driven motion generation, we train a text-conditioned diffusion model on the low-dimensional latent space obtained by VAE model (Section 3.3). (IV) Guided generation: Adopting training-free guided generation in inference enables multiple editing functions required in motion generation without additional training (Section 3.4). The outline of our text-to-motion generation model, MoLA, is shown in Figure 3.
3.1. Motion representation
@@
We build upon the motion representation used in (Guo et al., 2022; Plappert et al., 2016) and improve the representation to enable variable length generation and improve performance in our framework. The pose representation used in (Guo et al., 2022; Plappert et al., 2016) is expressed as $\bm{m}\in\mathbb{R}^{(4+12N_{j}+4)\times L}$, where $N_{j}$ is the number of joints. A motion can be represented as a sequence of this representation. The $i$-th $(1\leq i\leq L)$ pose is defined as follows:
(1) $\bm{m}^{i}=[\dot{r}_{a}^{i},\dot{r}_{x}^{i},\dot{r}_{z}^{i},r_{y}^{i},\bm{j}_{p}^{i},\bm{j}_{r}^{i},\bm{j}_{v}^{i},\bm{c}^{i}]$
@@
where $\dot{r}_{a}\in\mathbb{R}$ is root angular velocity along the $Y$-axis, $\dot{r}_{x}$ and $\dot{r}_{z}\in\mathbb{R}$ are root linear velocities on $XZ$-plane, $r_{y}$ is root height, $\bm{j}_{p}\in\mathbb{R}^{3(N_{j}-1)}$ is local joints positions, $\bm{j}_{r}\in\mathbb{R}^{6(N_{j}-1)}$ is rotations in root space, $\bm{j}_{v}\in\mathbb{R}^{3N_{j}}$ is velocities and $\bm{c}\in\mathbb{R}^{4}$ is binary foot-ground contact features by thresholding the heel and toe joint velocities. To achieve variable-length motion generation during stage 2 inference (explained in Section 3.3), we concatenate an activation variable $a^{i}\in\mathbb{R}$ for the $i$-th pose to Equation (1) as $[\bm{m}^{i},a^{i}]$. Additionally, to reduce the burden on the encoder in Section 3.2, we remove the substantially redundant information $\bm{j}_{v}$ and $\bm{c}$ from Equation (1) as $\tilde{\bm{m}}^{i}=[\dot{r}_{a}^{i},\dot{r}_{x}^{i},\dot{r}_{z}^{i},r_{y}^{i},\bm{j}_{p}^{i},\bm{c}^{i}]$, and then we set the vector for the encoder input in Section 3.2 as follows:
(2) $\bm{x}^{i}=[\tilde{\bm{m}}^{i},a^{i}]=[\dot{r}_{a}^{i},\dot{r}_{x}^{i},\dot{r}_{z}^{i},r_{y}^{i},\bm{j}_{p}^{i},\bm{j}_{r}^{i},a^{i}].$
@@
3.2. Stage 1: Continuous Motion Latent Representation with Adversarial Training
3.2.1. Learning a motion latent representation with VAE-GAN
@@
(3) $\mathcal{J}_{\text{VAE}}(\bm{\psi},\bm{\eta};\bm{x})=$ $~{}\mathbb{E}_{q_{\bm{\eta}}(\bm{z}|\bm{x})}[\|\bm{x}-g_{\bm{\psi}}(\bm{z})\|_{2}^{2}]+\lambda_{\text{reg}}D_{KL}(q_{\bm{\eta}}(\bm{z}|\bm{x})\|p(\bm{z})),$
where $\lambda_{\text{reg}}$ is a hyperparameter for balancing the two terms. The encoder $q_{\bm{\eta}}$ can be trained to produce a low-dimensional latent representation, and the decoder $g_{\bm{\psi}}$ can also be trained to accurately reconstruct the input motion sequences from the latent representations. Besides, we separate the decoder output into two components, denoted as $g_{\bm{\psi}}(\bm{z})=[\bm{m}^{\prime}(\bm{z}),\bm{a}^{\prime}(\bm{z})]$, where $\bm{m}^{\prime}(\bm{z})\in\mathbb{R}^{(N-1)\times L}$ and $\bm{a}^{\prime}(\bm{z})\in\mathbb{R}^{L}$ correspond to the original motion and proposed activation variable. We adopt a binary cross entropy (BCE) loss for the activation variable. The following $\mathcal{J}_{\text{MotionVAE}}$ is used as the loss function for the VAE part:
@@
(4) $\qquad+\lambda_{\text{act}}\mathcal{L}_{BCE}(\bm{a},\bm{a}^{\prime}(\bm{z})))]+\lambda_{\text{reg}}D_{KL}(q_{\bm{\eta}}(\bm{z}|\bm{x})\|p(\bm{z})),$
@@
To achieve high-quality generation, we need to push the limits of compression. Hence, we propose incorporating adversarial training into the motion VAE. More specifically, we introduce a discriminator, denoted as $f_{\bm{\phi}}:\mathbb{R}^{N\times L}\to\mathbb{R}$, that aims to distinguish real and reconstructed motions. The adversarial training is formulated as a two-player optimization between the VAE and the discriminator. The discriminator is trained by the maximization of $\mathbb{E}_{p(\bm{x})}\mathcal{L}_{\text{GAN}}(\bm{\phi};\bm{\psi},\bm{\eta},\bm{x})$ with respect to $\bm{\phi}$, where
@@
(5) $+\mathbb{E}_{q_{\bm{\eta}}(\bm{z}|\bm{x})}\left[\min\{0,-1-f_{\bm{\phi}}(g_{\bm{\psi}}(\bm{z}))\}\right].$
@@
(6) $\min_{\phi,\psi}\mathbb{E}_{p(x)}\left[\mathcal{J}_{\text{MotionVAE}}(\bm{\psi},\bm{\eta};\bm{x})+\lambda_{\text{adv}}\mathcal{J}_{\text{GAN}}(\bm{\psi},\bm{\eta};\bm{\phi},\bm{x})\right],$
@@
(7) $\mathcal{J}_{\text{GAN}}(\bm{\psi},\bm{\eta};\bm{\phi},\bm{x})=$ $~{}-\mathbb{E}_{q_{\bm{\eta}}(\bm{z}|\bm{x})}[f_{\bm{\phi}}(g_{\bm{\psi}}(\bm{z}))].$
@@
3.2.2. From GAN- to SAN-based discriminator

### v4

@@ Abstract
In text-to-motion generation, controllability as well as generation quality and speed has become increasingly critical. The controllability challenges include generating a motion of a length that matches the given textual description and editing the generated motions according to control signals, such as the start-end positions and the pelvis trajectory. In this paper, we propose MoLA, which provides fast, high-quality, variable-length motion generation and can also deal with multiple editing tasks in a single framework. Our approach revisits the motion representation used as inputs and outputs in the model, incorporating an activation variable to enable variable-length motion generation. Additionally, we integrate a variational autoencoder and a latent diffusion model, further enhanced through adversarial training, to achieve high-quality and fast generation. Moreover, we apply a training-free guided generation framework to achieve various editing tasks with motion control inputs. We quantitatively show the effectiveness of adversarial learning in text-to-motion generation, and demonstrate the applicability of our editing framework to multiple editing tasks in the motion domain. Our code is available at https://github.com/sony/MoLA.
@@
Figure 1: MoLA achieves fast and high-quality human motion generation given textual descriptions while enabling motion editing applications. With MoLA, we can deal with various types of motion editing tasks in a single framework.
1 Introduction
@@
Human motion synthesis from text is an emerging task with highly relevant applications in fields such as multimedia production and computer animation. For example, an animator might wish to create or edit a motion prototype to verify their artistic intent before time-consuming animation or motion capture commences, generate smooth in-between motion between two motion capture clips, or generate specific motion that follows a predefined trajectory. In Figure 1, we demonstrate results of our approach on such motion generation and editing tasks. To be useful in those real-world applications, a method has to excel in three domains: (1) motion quality, which encompasses both the general motion quality and the adherence to the textual description; (2) fast inference time; (3) efficient motion editing.
@@
Several recent works have attempted to address these desired properties: Methods based on vector quantization (VQ), such as T2M-GPT [44], MoMask [13], MMM [29], ParCo [49], and BAMM [28], achieve impressive generation quality by compressing human motion into discrete tokens and then sampling those tokens to synthesize motion. Diffusion-based methods in data space, such as MDM [37], provide impressive flexibility with regard to quality and motion editability. To further improve motion quality and inference time, latent-space based methods such as MLD [3] or MotionMamba [47] operate in a learned continuous latent space.
@@
However, no state-of-the-art method excels in all three domains necessary for real-world applications. For example, while latent-space based approaches such as VQ-based methods or MLD achieve impressive motion quality and fast inference time, they cannot edit a given motion sequence in a training-free manner. In contrast, data-space based methods such as MDM are able to edit motion in a training-free manner, which, however, comes at the cost of slow inference time and lower generation quality. Moreover, while these models excel at motion generation, they require users to manually specify the motion length instead of automatically determining it from the textual input. This often necessitates length estimation or iterative adjustment, which limits their flexibility. For instance, MoMask [13] uses a length estimator during inference to predict motion length based on text input. However, inaccurate predictions may result in significant motion drift [40].
@@
To close this gap, we propose MoLA, Motion Generation and Editing with Latent Diffusion Enhanced by Adversarial Training. We revisit the representation of motion features used as inputs and outputs of the model and introduce an activation variable that characterizes the length of the motion. We utilize a variational auto-encoder (VAE) [20] for its continuous latent space, which allows training-free editing. This is in contrast to discrete latent spaces that do not allow for training-free motion editing. We also enhance the VAE training with adversarial learning, which has been shown to work well in the image [8, 16, 32] and audio [23] domains. We empirically show that our model achieves variable-length motion generation aligned with textual descriptions and high generation performance on a commonly used dataset [12]. We also demonstrate that training-free guided diffusion can enable multiple motion editing tasks such as path-following, in-betweening, and upper body editing. These experiments also highlight the significant improvement in speed and performance of our model compared to the performance of existing training-free motion editing models (see Figure 2).
@@
The main contributions of this paper are threefold. First, we introduce an activation variable into the motion representation and show that our method can perform variable-length motion generation conditioned on text. Second, we propose a new continuous latent-based motion generation model that introduces adversarial training into the motion VAE and quantitatively show that it significantly pushes the limits of existing continuous-based methods. Finally, we demonstrate that our model not only shows high generation performance but also can deal with various types of motion editing tasks in a training-free manner.
@@
Figure 2: Comparison of inference cost, generation performance, and editability for text-to-motion methods on HumanML3D dataset (All tests are performed on NVIDIA A100 GPU). $\bullet$ means a method that can edit motion in a training-free manner, and $\times$ means a method that cannot edit motion in a training-free manner. The pink arrow in the figure indicates that MoLA significantly extends the performance boundaries (in terms of generation quality and speed) of methods categorized as enabling training-free editing.
2 Related Work
2.1 Motion Generation
@@
Text-to-motion generation technology has made rapid progress with the diffusion models and VQ-based models. MDM [37] and MotionDiffuse [46] adopt diffusion models for motion generation, which leads to better performance in terms of generation quality. However, these methods directly apply diffusion processes to raw motion sequences, thus resulting in slow generation. To address this issue, MLD [3] adopts a diffusion model in a low-dimensional latent space provided by a VAE trained on motion data, drawing inspiration from latent diffusion models [32]. VQ-based models have been studied as well, inspired by the success of VQ-based models in image generation [10, 42, 2]. In this approach, a VQ-VAE model is first trained on motion data to acquire discrete motion representations, and deep generative models are then applied to generate sequences of discrete representations (also called tokens). T2M-GPT [44], AttT2M [48], MotionGPT [17] and ParCo [49] utilize autoregressive (AR) models to generate motion tokens. However, AR models are slow in inference because motion tokens are generated sequentially. To address this issue, M2DM [21] and DiverseMotion [24] apply discrete diffusion models to motion tokens in a latent space, whereas MMM [29], MoMask [13] and BAMM [28] adopt a mask prediction model.
@@
Figure 3: The overall framework of MoLA. Stage 1: A motion VAE enhanced by adversarial training learns a low-dimensional latent representation of diverse motion sequences. Stage 2: A text-conditioned latent diffusion model leverages this representation for fast and high-quality text-to-motion generation. Guided generation: During inference, a gradient-based method minimizes a loss function $\mathfrak{L}_{\text{Motion}}$ for each desired editing task, enabling multiple motion editing tasks within a unified framework.
2.2 Motion Editing
@@
Motion editing has attracted much research interest as well. MDM [37] demonstrated upper body editing and motion in-betweening by applying diffusion inpainting to motion data in both the spatial and temporal domains. LGD [34] demonstrated path-following motion generation with a guided diffusion that utilizes multiple samples from a suitable distribution to reduce bias. GMD [18] guides the position of the root joint to control motion trajectories. OmniControl [41] controls any joints at any time by guiding a pre-trained motion diffusion model with an analytic function. DNO [19] optimizes the diffusion latent noise of a pre-trained text-to-motion model with user-provided criteria in the motion space, achieving multiple editing tasks. However, these methods employ data-space diffusion models similar to MDM and have not been demonstrated with a latent diffusion model. MMM [29] demonstrated motion editing by placing masked tokens where editing is needed and applying a mask prediction framework, and MotionLCM [6] proposed a fast controllable motion generation framework by introducing latent consistency distillation and the motion ControlNet [45] manipulation in the latent space.
3 Method
@@
The goal of this study is to develop a framework for fast and high-quality text-guided motion generation and to deal with multiple control tasks in a training-free way. To achieve this, we propose the following training and inference tricks (I)-(IV): (I) We review the representation of motion features to achieve improved accuracy and variable length motion generation. We select the necessary features to reduce the burden on our model’s VAE encoder and add an activation variable to the representation to determine the motion length (Section 3.1). (II) We train a motion VAE enhanced by adversarial training to achieve high quality generation performance (Section 3.2). (III) To reduce computational complexity while simultaneously enabling high-quality text-driven motion generation, we train a text-conditioned diffusion model on the low-dimensional latent space obtained by VAE model (Section 3.3). (IV) Guided generation: Adopting training-free guided generation in inference enables multiple editing functions required in motion generation without additional training (Section 3.4). The outline of our text-to-motion generation model, MoLA, is shown in Figure 3.
3.1 Motion representation
@@
We build upon the motion representation used in [12, 30] and improve the representation to enable variable length generation and improve performance in our framework. The pose representation used in [12, 30] is expressed as $\bm{m}\in\mathbb{R}^{(4+12N_{j}+4)\times L}$, where $N_{j}$ is the number of joints. A motion can be represented as a sequence of this representation. The $i$-th $(1\leq i\leq L)$ pose is defined as follows:
$\bm{m}^{i}=[\dot{r}_{a}^{i},\dot{r}_{x}^{i},\dot{r}_{z}^{i},r_{y}^{i},(\bm{j}_{p}^{i})^{\top},(\bm{j}_{r}^{i})^{\top},(\bm{j}_{v}^{i})^{\top},(\bm{c}^{i})^{\top}]^{\top}$ (1)
@@
where $\dot{r}_{a}^{i}\in\mathbb{R}$ is root angular velocity along the $Y$-axis, $\dot{r}_{x}^{i}$ and $\dot{r}_{z}^{i}\in\mathbb{R}$ are root linear velocities on $XZ$-plane, $r_{y}^{i}$ is root height, $\bm{j}_{p}^{i}\in\mathbb{R}^{3(N_{j}-1)}$ is local joints positions, $\bm{j}_{r}^{i}\in\mathbb{R}^{6(N_{j}-1)}$ is rotations in root space, $\bm{j}_{v}^{i}\in\mathbb{R}^{3N_{j}}$ is velocities and $\bm{c}^{i}\in\mathbb{R}^{4}$ is binary foot-ground contact features by thresholding the heel and toe joint velocities. To achieve variable-length motion generation during stage 2 inference (explained in Section 3.3), we concatenate an activation variable $a^{i}\in\mathbb{R}$ for the $i$-th pose to Equation (1) as $[(\bm{m}^{i})^{\top},a^{i}]^{\top}$. Additionally, to reduce the burden on the encoder in Section 3.2, we remove the substantially redundant information $\bm{j}_{v}$ and $\bm{c}$ from Equation (1) as $\tilde{\bm{m}}^{i}=[\dot{r}_{a}^{i},\dot{r}_{x}^{i},\dot{r}_{z}^{i},r_{y}^{i},(\bm{j}_{p}^{i})^{\top},(\bm{c}^{i})^{\top}]^{\top}$, and then we set the vector for the encoder input in Section 3.2 as follows:
$\bm{x}^{i}=[(\tilde{\bm{m}}^{i})^{\top},a^{i}]^{\top}=[\dot{r}_{a}^{i},\dot{r}_{x}^{i},\dot{r}_{z}^{i},r_{y}^{i},(\bm{j}_{p}^{i})^{\top},(\bm{j}_{r}^{i})^{\top},a^{i}]^{\top}.$ (2)
@@
3.2 Stage 1: Continuous Motion Latent Representation with Adversarial Training
3.2.1 Learning a motion latent representation with VAE-GAN
@@
$\mathcal{J}_{\text{VAE}}(\bm{\psi},\bm{\eta};\bm{x})=$ $~{}\mathbb{E}_{q_{\bm{\eta}}(\bm{z}|\bm{x})}[\|\bm{x}-g_{\bm{\psi}}(\bm{z})\|_{2}^{2}]$
$\qquad+\lambda_{\text{reg}}D_{KL}(q_{\bm{\eta}}(\bm{z}|\bm{x})\|p(\bm{z})),$ (3)
where $\lambda_{\text{reg}}$ is a hyperparameter for balancing the two terms. The encoder $q_{\bm{\eta}}$ can be trained to produce a low-dimensional latent representation, and the decoder $g_{\bm{\psi}}$ can also be trained to accurately reconstruct the input motion sequences from the latent representations. Besides, we separate the decoder output into two components, denoted as $g_{\bm{\psi}}(\bm{z})=[(\bm{m}^{\prime}(\bm{z}))^{\top},\bm{a}^{\prime}(\bm{z})]^{\top}$, where $\bm{m}^{\prime}(\bm{z})\in\mathbb{R}^{(N-1)\times L}$ and $\bm{a}^{\prime}(\bm{z})\in\mathbb{R}^{L}$ correspond to the original motion and proposed activation variable. We adopt a binary cross entropy (BCE) loss for the activation variable. The following $\mathcal{J}_{\text{MotionVAE}}$ is used as the loss function for the VAE part:
@@
$\qquad+\lambda_{\text{act}}\mathcal{L}_{BCE}(\bm{a},\bm{a}^{\prime}(\bm{z})))]+\lambda_{\text{reg}}D_{KL}(q_{\bm{\eta}}(\bm{z}|\bm{x})\|p(\bm{z})),$ (4)
@@
To achieve high-quality generation, we need to push the limits of compression. Hence, we propose incorporating adversarial training into the motion VAE (c.f. [32, 23]). More specifically, we introduce a discriminator, denoted as $f_{\bm{\phi}}:\mathbb{R}^{N\times L}\to\mathbb{R}$, that aims to distinguish real and reconstructed motions. The adversarial training is formulated as a two-player optimization between the VAE and the discriminator. The discriminator is trained by the maximization of $\mathbb{E}_{p(\bm{x})}\mathcal{L}_{\text{GAN}}(\bm{\phi};\bm{\psi},\bm{\eta},\bm{x})$ with respect to $\bm{\phi}$, where
@@
$\qquad\qquad\quad+\mathbb{E}_{q_{\bm{\eta}}(\bm{z}|\bm{x})}\left[\min\{0,-1-f_{\bm{\phi}}(g_{\bm{\psi}}(\bm{z}))\}\right].$ (5)
@@
$\min_{\phi,\psi}\mathbb{E}_{p(x)}\left[\mathcal{J}_{\text{MotionVAE}}(\bm{\psi},\bm{\eta};\bm{x})+\lambda_{\text{adv}}\mathcal{J}_{\text{GAN}}(\bm{\psi},\bm{\eta};\bm{\phi},\bm{x})\right],$ (6)
@@
$\mathcal{J}_{\text{GAN}}(\bm{\psi},\bm{\eta};\bm{\phi},\bm{x})=$ $~{}-\mathbb{E}_{q_{\bm{\eta}}(\bm{z}|\bm{x})}[f_{\bm{\phi}}(g_{\bm{\psi}}(\bm{z}))].$ (7)
@@
3.2.2 From GAN- to SAN-based discriminator